OpenStudy (anonymous):

please prove this statement sir if a and b are prime to one another,then nos;1a,2a,3a....(b-1)a leave different remaining when divided by b

OpenStudy (anonymous):

try a proof by contradiction is my first guess. suppose \(a_i\) and \(a_j\) leave the same remainder when divide by \(b\) , call that remainder \(r\) and see if you can show that that means \(gcd(a, b)\neq 1\)

OpenStudy (anonymous):

rather i meant \(ia\) and \(ja\) leave the same remainder ignore the subscripts, they were stupid

OpenStudy (anonymous):

hint, consider \(ia-ja=(i-j)a\)

OpenStudy (anonymous):

sir its proof is if possible let ra and sa leave the same remainders when divided by b,then ra-sa=m(b) (r-s)a=m(b) since a is prime to be a not equal to m(b)such tha r-s=m(b) which is contradiction where r,s<b such that r-s<b so our assemption is wrong so the remainders are different pls help me sir for prove this with adding numbers